OpenStudy (anonymous):

Simplify the expression \[\sqrt{ #4 } \] divided by (5 + 2i) - (3 - 4i)

OpenStudy (mertsj):

\[\frac{\sqrt{4}}{5+2i-(3-4i)}\]

OpenStudy (anonymous):

multiply by conjugate of the denominator, which is 2+2i

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yes, mertsj

OpenStudy (anonymous):

i put 5 + 2i in parentheses, but it doesn't really matter either way

OpenStudy (anonymous):

okay myoko

OpenStudy (anonymous):

so multiply bouth nominator and denominator by 2+2i and see what you get

OpenStudy (anonymous):

numerator would be 3

OpenStudy (anonymous):

\[\frac{\sqrt4(2+2i)}{(2-2i)(2+2i)}=\frac{\sqrt4(2+2i)}{8}=+-\frac{2+2i}{4}\]

OpenStudy (anonymous):

+-(1/2+1/2i)

OpenStudy (sirm3d):

the conjugate of the denominator is 2-6i

OpenStudy (anonymous):

right. Ty @sirm3d

OpenStudy (anonymous):

just change 2+2i for 2-6i and perfomr the operations